The bike is moving at a speed of 18 km / h. The wheel of the bicycle has a radius of 25 cm. How many revolutions does it make in 15s.


1
Expert's answer
2021-03-04T11:51:41-0500

Let's change the frame of reference, and consider the situation from the wheel's center perspective. In this frame of reference all points of the wheel move with speed "v = 18km\/h = 5m\/s" in the circle. Thus, in "t = 15s" a point on the wheel cover the following distance:


"d = vt = 75m"


This distance correspondes to the following amount of revolution:


"N = \dfrac{d}{2\pi R}"

where "2\pi R" is the length of the wheel ("R = 25cm = 0.25m"). Finally, obtain:


"N = \dfrac{75}{2\pi\cdot 0.25} \approx 47.7"

Answer. 47.7.